# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v3

@@ 4 Conclusions
In this work, we presented JWST/MIRI MRS observations of the active galaxy NGC 1365 as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. We analyzed the spatially resolved gas properties and compared them with optical and mm observations from MUSE and ALMA, respectively. We traced both the ionized atomic and the warm and cold molecular gas phases in the circumnuclear region ( 0.90.9 kpc2) of the NGC 1365 galaxy, identifying more than 40 mid-IR emission lines. Our main results are the following:
@@
We find that the mid-IR, optical, and mm emission lines consistently trace different phases of the gas with coherent internal kinematics within each phase (see Fig. 3 and Section 3.1). The cold molecular gas traced by the CO(3-2) transition shares the same ordered kinematics of the stellar galaxy disk (Venturi et al., 2018), with no evidence for outflowing material. The same conclusion holds for the warm molecular gas traced by the H2 pure-rotational transitions, with the exception of a peculiar high- unresolved region located southwards of the nucleus.
@@
The ionized gas lines can be separated in two main groups which show opposed morphologies based on their ionization potential (IP). The high-IP ( 54 eV) emission lines trace the bipolar outflow, showing a velocity gradient perpendicular to the disk major axis, consistent with the [O III] kinematics. The velocity dispersion map highlights a donut-shaped structure associated with the approaching side of the outflow which extends up to 2.5″ ( 240 pc). The low-IP ( 25 eV) emission line kinematics resembles the stellar galaxy disk ordered motions, similar to the H2, with clumpy flux peaks in the circumnuclear star-forming ring. Finally, medium-IP lines show intermediate properties, with a kinematics similar to low-IP species but exhibiting the donut-shaped structure typical of the high-IP lines in the velocity dispersion map.
@@
[O III] and [Ne V] velocity channel maps reveal that both outflowing species trace the same kinematic structures, with the [Ne V] appearing more collimated on the outflow axis, as expected from their different IPs (Fig. 4 and Section 3.2). The [Ne V] emission also reveals the NW receding counterpart of the ionization cone, which is undetected in [O III] due to large extinction caused by the galaxy disk at optical wavelengths.
@@
The mid-IR emission-line ratios indicate a composite excitation scenario, with contribution from both AGN photoionization and shocks (Fig. 5 and Section 3.3). The AGN ionization seems to extend along the disk major axis, while in the other direction we observe shock-dominated zones, suggesting an interaction between the outflow and the disk material. Notably, by penetrating the dust in the central regions of the galaxy, these mid-IR diagnostic diagrams reveal for the first time the role of the AGN in ionizing the gas in the nuclear region, differently from the optical diagnostic diagrams that mostly reveal excitation due to SF.
@@
By exploiting the innovative photoionization modeling code HOMERUN, we derived the physical properties of the ionized outflowing gas, providing a self-consistent and physically motivated estimate of the outflow mass (Section 3.5). We obtain = (109 2) 103 and = (88 9) 103 . These estimates differ significantly from those derived with the standard method, which are 40 times lower ([Ne V]) and 2 times higher ([O III]) than those obtained from HOMERUN. This discrepancy reflects the strong assumptions in the classical approach regarding ionization state, average density, and extinction correction.
@@
This work highlights the critical need for self-consistent, multi-line photoionization and kinematical modeling to properly capture the complex structure and energetics of AGN-driven outflows - a key step toward understanding their impact on galaxy evolution. We fully exploited the synergy between cutting-edge IFU facilities, JWST/MIRI and VLT/MUSE, to model the ionized gas phase properties across a broad ionization range (from 10 eV to 130 eV), using a total of 60 emission lines. For the first time, we combined detailed 3D kinematic modeling with and photoionization fitting with HOMERUN to derive the intrinsic structure and physical conditions of the outflow in a fully self-consistent framework. This unprecedented approach enabled a coherent and physically motivated characterization of the outflow’s spatial structure, ionization stratification, and energetics. Our results show that classical methods, based on simplified assumptions and a limited set of tracers, significantly underestimate the outflow energetics - and hence the strength of AGN feedback.
@@ Acknowledgements.
We are greatful to S. Charlot for kindly providing the predictions for the SF line-emission models. EB, FB, and GC acknowledge financial support from INAF under the Large Grant 2022 “The metal circle: a new sharp view of the baryon cycle up to Cosmic Dawn with the latest generation IFU facilities” and the GO grant 2024 “A JWST/MIRI MIRACLE: Mid-IR Activity of Circumnuclear Line Emission”. IL, FB, and AM acknowledge support from PRIN-MUR project “PROMETEUS” financed by the European Union - Next Generation EU, Mission 4 Component 1 CUP B53D23004750006. EB acknowledges INAF funding through the “Ricerca Fondamentale 2024” program (mini-grant 1.05.24.07.01). IL and AM acknowledges support from PRIN-MUR project “PROMETEUS” financed by the European Union - Next Generation EU, Mission 4 Component 1 CUP B53D23004750006. AM, MG, IL and CM acknowledge INAF funding through the “Ricerca Fondamentale 2023” program (mini-grant 1.05.23.04.01). FS acknowledges financial support from the PRIN MUR 2022 2022TKPB2P - BIG-z, Ricerca Fondamentale INAF 2023 Data Analysis grant 1.05.23.03.04 “ARCHIE ARchive Cosmic HI & ISM Evolution”, Ricerca Fondamentale INAF 2024 under project 1.05.24.07.01 MINI-GRANTS RSN1 ”ECHOS”, Bando Finanziamento ASI CI-UCO-DSR-2022-43 project “IBISCO: feedback and obscuration in local AGN”. GS acknowledges the project PRIN-MUR 2020 MUR BEYOND-2p (“Astrochemistry beyond the second period elements”, Prot. 2020AFB3FX), the PRIN MUR 2022 FOSSILS (“Chemical origins: linking the fossil composition of the Solar System with the chemistry of protoplanetary disks”, Prot. 2022JC2Y93), the project ASI-Astrobiologia 2023 MIGLIORA (“Modeling Chemical Complexity”, F83C23000800005), the INAF-GO 2023 fundings PROTOSKA (“Exploiting ALMA data to study planet forming disks: preparing the advent of SKA”, C13C23000770005) and the INAF-Minigrant 2023 TRIESTE (“TRacing the chemIcal hEritage of our originS: from proTostars to planEts”; PI: G. Sabatini). SC and GV acknowledge support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. JF acknowledges financial support from CONAHCyT, project number CF-2023-G100, and UNAM-DGAPA-PAPIIT IN111620 grant, Mexico. AVG acknowledges support from the Spanish grant PID2022-138560NB-I00, funded by MCIN/AEI/10.13039/501100011033/FEDER, EU. MM is thankful for support from the European Space Agency (ESA). EH gratefully acknowledges the hospitality of the IAC, where part of this work was carried out during a long research visit. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The specific observations analyzed can be accessed via doi: https://doi.org/10.17909/b4w1-hk44. These observations are associated with program 6138. The authors acknowledge the team led by coPIs C. Marconcini and A. Feltre for developing their observing program with a zero-exclusive-access period. This paper makes use of the following ALMA data: ADS/JAO.ALMA2016.1.00296.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.References
- Alarie & Morisset (2019) Alarie, A. & Morisset, C. 2019, Rev. Mexicana Astron. Astrofis., 55, 377
- Alonso-Herrero et al. (2012) Alonso-Herrero, A., Sánchez-Portal, M., Ramos Almeida, C., et al. 2012, MNRAS, 425, 311
- Anand et al. (2021) Anand, G. S., Lee, J. C., Van Dyk, S. D., et al. 2021, MNRAS, 501, 3621
- Argyriou et al. (2023) Argyriou, I., Glasse, A., Law, D. R., et al. 2023, A&A, 675, A111
- Argyriou et al. (2020) Argyriou, I., Wells, M., Glasse, A., et al. 2020, A&A, 641, A150
- Asplund et al. (2021) Asplund, M., Amarsi, A. M., & Grevesse, N. 2021, A&A, 653, A141
- Asplund et al. (2009) Asplund, M., Grevesse, N., Sauval, A. J., & Scott, P. 2009, ARA&A, 47, 481
- Audibert et al. (2023) Audibert, A., Ramos Almeida, C., García-Burillo, S., et al. 2023, A&A, 671, L12
@@
- Bacon et al. (2010) Bacon, R., Accardo, M., Adjali, L., et al. 2010, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 7735, Ground-based and Airborne Instrumentation for Astronomy III, ed. I. S. McLean, S. K. Ramsay, & H. Takami, 773508
- Baldwin et al. (1981) Baldwin, J. A., Phillips, M. M., & Terlevich, R. 1981, PASP, 93, 5
- Belli et al. (2024) Belli, S., Park, M., Davies, R. L., et al. 2024, Nature, 630, 54
- Boroson & Green (1992) Boroson, T. A. & Green, R. F. 1992, ApJS, 80, 109
- Braito et al. (2014) Braito, V., Reeves, J. N., Gofford, J., et al. 2014, ApJ, 795, 87
- Bureau et al. (1996) Bureau, M., Mould, J. R., & Staveley-Smith, L. 1996, ApJ, 463, 60
- Bushouse et al. (2022) Bushouse, H., Eisenhamer, J., Dencheva, N., et al. 2022, spacetelescope/jwst: JWST 1.6.2
- Byrne et al. (2022) Byrne, C. M., Stanway, E. R., Eldridge, J. J., McSwiney, L., & Townsend, O. T. 2022, MNRAS, 512, 5329
- Cano-Díaz et al. (2012) Cano-Díaz, M., Maiolino, R., Marconi, A., et al. 2012, A&A, 537, L8
- Cappellari & Copin (2003) Cappellari, M. & Copin, Y. 2003, MNRAS, 342, 345
- Cappellari & Emsellem (2004) Cappellari, M. & Emsellem, E. 2004, PASP, 116, 138
- Carniani et al. (2015) Carniani, S., Marconi, A., Maiolino, R., et al. 2015, A&A, 580, A102
- CASA Team et al. (2022) CASA Team, Bean, B., Bhatnagar, S., et al. 2022, PASP, 134, 114501
- Ceci et al. (2025) Ceci, M., Cresci, G., Arribas, S., et al. 2025, A&A, 695, A116
- Cicone et al. (2018) Cicone, C., Brusa, M., Ramos Almeida, C., et al. 2018, Nature Astronomy, 2, 176
- Combes (2017) Combes, F. 2017, Frontiers in Astronomy and Space Sciences, 4, 10
- Combes et al. (2019) Combes, F., García-Burillo, S., Audibert, A., et al. 2019, A&A, 623, A79
- Cresci & Maiolino (2018) Cresci, G. & Maiolino, R. 2018, Nature Astronomy, 2, 179
- Cresci et al. (2015) Cresci, G., Marconi, A., Zibetti, S., et al. 2015, A&A, 582, A63
- Cresci et al. (2023) Cresci, G., Tozzi, G., Perna, M., et al. 2023, A&A, 672, A128
- Davies et al. (2024) Davies, R., Shimizu, T., Pereira-Santaella, M., et al. 2024, A&A, 689, A263

### v4

@@ 4 Conclusions
In this work, we have presented JWST/MIRI MRS observations of the active galaxy NGC 1365 as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. We analyzed the spatially resolved gas properties and compared them with optical and millimeter observations from MUSE and ALMA, respectively. We traced both the ionized atomic and the warm and cold molecular gas phases in the circumnuclear region ( 0.90.9 kpc2) of the NGC 1365 galaxy, identifying more than 40 mid-IR emission lines. Our main results are the following:
@@
We find that the mid-IR, optical, and millimeter emission lines consistently trace different phases of the gas with coherent internal kinematics within each phase (see Fig. 3 and Section 3.1). The cold molecular gas traced by the CO(3-2) transition shares the same ordered kinematics of the stellar galaxy disk (Venturi et al., 2018), with no evidence of outflowing material. The same conclusion holds for the warm molecular gas traced by the H2 pure-rotational transitions, with the exception of a peculiar high- unresolved region located southward of the nucleus.
@@
The ionized gas lines can be separated in two main groups that show opposed morphologies based on their IP. The high-IP ( 54 eV) emission lines trace the bipolar outflow, showing a velocity gradient perpendicular to the disk major axis, consistent with the [O III] kinematics. The velocity dispersion map highlights a donut-shaped structure associated with the approaching side of the outflow, which extends up to 2.5″ ( 240 pc). The low-IP ( 25 eV) emission line kinematics resembles the stellar galaxy disk ordered motions, similar to the H2, with clumpy flux peaks in the circumnuclear star-forming ring. Finally, medium-IP lines show intermediate properties, with a kinematics similar to low-IP species but exhibiting the donut-shaped structure typical of the high-IP lines in the velocity dispersion map.
@@
[O III] and [Ne V] velocity channel maps reveal that both outflowing species trace the same kinematic structures, with the [Ne V] appearing more collimated on the outflow axis, as was expected from their different IPs (Fig. 4 and Section 3.2). The [Ne V] emission also reveals the NW receding counterpart of the ionization cone, which is undetected in [O III] due to large extinction caused by the galaxy disk at optical wavelengths.
@@
The mid-IR emission-line ratios indicate a composite excitation scenario, with contributions from both AGN photoionization and shocks (Fig. 5 and Section 3.3). The AGN ionization seems to extend along the disk major axis, while in the other direction we observe shock-dominated zones, suggesting an interaction between the outflow and the disk material. Notably, by penetrating the dust in the central regions of the galaxy, these mid-IR diagnostic diagrams reveal for the first time the role of the AGN in ionizing the gas in the nuclear region, differently from the optical diagnostic diagrams that mostly reveal excitation due to SF.
@@
By exploiting the innovative photoionization modeling code HOMERUN, we derived the physical properties of the ionized outflowing gas, providing a self-consistent and physically motivated estimate of the outflow mass (Section 3.5). We obtained = (109 2) 103 and = (88 9) 103 . These estimates differ significantly from those derived with the standard method, which are 40 times lower ([Ne V]) and 2 times higher ([O III]) than those obtained from HOMERUN. This discrepancy reflects the strong assumptions in the classical approach regarding ionization state, average density, and extinction correction.
@@
This work highlights the critical need for self-consistent, multiline photoionization and kinematical modeling to properly capture the complex structure and energetics of AGN-driven outflows – a key step toward understanding their impact on galaxy evolution. We fully exploited the synergy between cutting-edge IFU facilities, JWST/MIRI and VLT/MUSE, to model the ionized gas phase properties across a broad ionization range (from 10 eV to 130 eV), using a total of 60 emission lines. For the first time, we combined detailed 3D kinematic modeling with and photoionization fitting with HOMERUN to derive the intrinsic structure and physical conditions of the outflow in a fully self-consistent framework. This unprecedented approach enabled a coherent and physically motivated characterization of the outflow’s spatial structure, ionization stratification, and energetics. Our results show that classical methods, based on simplified assumptions and a limited set of tracers, significantly underestimate the outflow energetics – and hence the strength of AGN feedback.
5 Data availability
The moment maps of the other emission lines listed in Table LABEL:tab:_table_list_emission_lines are shown in the online material. The reduced MIRI IFU datacubes used in this work are available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr(xxxxxxxx) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/xxx/xxx.
@@ Acknowledgements.
We are greatful to S. Charlot for kindly providing the predictions for the SF line-emission models. EB, FB, and GC acknowledge financial support from INAF under the Large Grant 2022 “The metal circle: a new sharp view of the baryon cycle up to Cosmic Dawn with the latest generation IFU facilities” and the GO grant 2024 “A JWST/MIRI MIRACLE: Mid-IR Activity of Circumnuclear Line Emission”. IL, FB, and AM acknowledge support from PRIN-MUR project “PROMETEUS” financed by the European Union - Next Generation EU, Mission 4 Component 1 CUP B53D23004750006 and C53D2300080006. EB acknowledges INAF funding through the “Ricerca Fondamentale 2024” program (mini-grant 1.05.24.07.01). AM, MG, IL and CM acknowledge INAF funding through the “Ricerca Fondamentale 2023” program (mini-grant 1.05.23.04.01). FS acknowledges financial support from the PRIN MUR 2022 2022TKPB2P - BIG-z, Ricerca Fondamentale INAF 2023 Data Analysis grant 1.05.23.03.04 “ARCHIE ARchive Cosmic HI & ISM Evolution”, Ricerca Fondamentale INAF 2024 under project 1.05.24.07.01 MINI-GRANTS RSN1 ”ECHOS”, Bando Finanziamento ASI CI-UCO-DSR-2022-43, CUP C93C25004260005, project “IBISCO: feedback and obscuration in local AGN”. G.S. acknowledges the project ASI-Astrobiologia 2023 MIGLIORA (“Modeling Chemical Complexity”, F83C23000800005), the INAF-GO 2023 fundings PROTOSKA (“Exploiting ALMA data to study planet forming disks: preparing the advent of SKA”, C13C23000770005), the INAF Minigrant 2023 TRIESTE (“TRacing the chemIcal hEritage of our originS: from proTostars to planEts”; PI: G. Sabatini) and financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU-Project Title 2022JC2Y93 Chemical Origins: linking the fossil composition of the Solar System with the chemistry of protoplanetary disks – CUP J53D23001600006 – Grant Assignment Decree No. 962 adopted on 30.06.2023 by the Italian Ministry of Ministry of University and Research (MUR). SC and GV acknowledge support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. JF acknowledges financial support from CONAHCyT, project number CF-2023-G100, and UNAM-DGAPA-PAPIIT IN111620 grant, Mexico. AVG acknowledges support from the Spanish grant PID2022-138560NB-I00, funded by MCIN/AEI/10.13039/501100011033/FEDER, EU. MM is thankful for support from the European Space Agency (ESA). EH gratefully acknowledges the hospitality of the IAC, where part of this work was carried out during a long research visit. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The specific observations analyzed can be accessed via doi: https://doi.org/10.17909/b4w1-hk44. These observations are associated with program 6138. The authors acknowledge the team led by coPIs C. Marconcini and A. Feltre for developing their observing program with a zero-exclusive-access period. This paper makes use of the following ALMA data: ADS/JAO.ALMA2016.1.00296.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.References
- Extensive Online Shock Model Database. Rev. Mexicana Astron. Astrofis. 55, pp. 377–394. External Links: Document, 1908.08579, ADS entry Cited by: §3.3.
- Probing the nuclear and circumnuclear activity of NGC 1365 in the infrared. MNRAS 425 (1), pp. 311–324. External Links: Document, 1206.2108, ADS entry Cited by: §1.
- Distances to PHANGS galaxies: New tip of the red giant branch measurements and adopted distances. MNRAS 501 (3), pp. 3621–3639. External Links: Document, 2012.00757, ADS entry Cited by: §1.
- JWST MIRI flight performance: The Medium-Resolution Spectrometer. A&A 675, pp. A111. External Links: Document, 2303.13469, ADS entry Cited by: §A.1, §2.
- The nature of point source fringes in mid-infrared spectra acquired with the James Webb Space Telescope. A&A 641, pp. A150. External Links: Document, 2007.16143, ADS entry Cited by: §A.1.
- The chemical make-up of the Sun: A 2020 vision. A&A 653, pp. A141. External Links: Document, 2105.01661, ADS entry Cited by: §3.5.2.
- The Chemical Composition of the Sun. ARA&A 47 (1), pp. 481–522. External Links: Document, 0909.0948, ADS entry Cited by: Appendix G, Appendix G.
- Jet-induced molecular gas excitation and turbulence in the Teacup. A&A 671, pp. L12. External Links: Document, 2302.13884, ADS entry Cited by: §1.
@@
- The MUSE second-generation VLT instrument. In Ground-based and Airborne Instrumentation for Astronomy III, I. S. McLean, S. K. Ramsay, and H. Takami (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 7735, pp. 773508. External Links: Document, 2211.16795, ADS entry Cited by: §1.
- Classification parameters for the emission-line spectra of extragalactic objects.. PASP 93, pp. 5–19. External Links: Document, ADS entry Cited by: §1.
- Star formation shut down by multiphase gas outflow in a galaxy at a redshift of 2.45. Nature 630 (8015), pp. 54–58. External Links: Document, 2308.05795, ADS entry Cited by: §1.
- The Emission-Line Properties of Low-Redshift Quasi-stellar Objects. ApJS 80, pp. 109. External Links: Document, ADS entry Cited by: §C.2.
- NGC 1365: A Low Column Density State Unveiling a Low Ionization Disk Wind. ApJ 795 (1), pp. 87. External Links: Document, 1409.3390, ADS entry Cited by: §1.
- A New I-Band Tully-Fisher Relation for the Fornax Cluster: Implication for the Fornax Distance and Local Supercluster Velocity Field. ApJ 463, pp. 60. External Links: Document, ADS entry Cited by: §1.
- spacetelescope/jwst: JWST 1.6.2 External Links: Document, ADS entry Cited by: §A.1.
- The dependence of theoretical synthetic spectra on -enhancement in young, binary stellar populations. MNRAS 512 (4), pp. 5329–5338. External Links: Document, 2203.13275, ADS entry Cited by: §3.5.1.
- Observational evidence of quasar feedback quenching star formation at high redshift. A&A 537, pp. L8. External Links: Document, 1112.3071, ADS entry Cited by: §1, §3.5.4.
- Adaptive spatial binning of integral-field spectroscopic data using Voronoi tessellations. MNRAS 342 (2), pp. 345–354. External Links: Document, astro-ph/0302262, ADS entry Cited by: §C.2.
- Parametric Recovery of Line-of-Sight Velocity Distributions from Absorption-Line Spectra of Galaxies via Penalized Likelihood. PASP 116 (816), pp. 138–147. External Links: Document, astro-ph/0312201, ADS entry Cited by: §C.2, §3.5.2.
- Ionised outflows in z ~2.4 quasar host galaxies. A&A 580, pp. A102. External Links: Document, 1506.03096, ADS entry Cited by: Appendix G, §3.5.4, §3.6.
- CASA, the Common Astronomy Software Applications for Radio Astronomy. PASP 134 (1041), pp. 114501. External Links: Document, 2210.02276, ADS entry Cited by: §A.3.
- The JWST/NIRSpec view of the nuclear region in the prototypical merging galaxy NGC 6240. A&A 695, pp. A116. External Links: Document, 2412.14685, ADS entry Cited by: §C.1, §1, §3.4.
- The largely unconstrained multiphase nature of outflows in AGN host galaxies. Nature Astronomy 2, pp. 176–178. External Links: Document, 1802.10308, ADS entry Cited by: §1.
- ALMA observations of molecular tori around massive black holes. A&A 623, pp. A79. External Links: Document, 1811.00984, ADS entry Cited by: §2, §3.1.
- AGN feedback and its quenching efficiency. Frontiers in Astronomy and Space Sciences 4, pp. 10. External Links: Document, 1707.09621, ADS entry Cited by: §1.
- The MAGNUM survey: positive feedback in the nuclear region of NGC 5643 suggested by MUSE. A&A 582, pp. A63. External Links: Document, 1508.04464, ADS entry Cited by: §1, §2.
- Bubbles and outflows: The novel JWST/NIRSpec view of the z = 1.59 obscured quasar XID2028. A&A 672, pp. A128. External Links: Document, 2301.11060, ADS entry Cited by: §1, §3.6.
- Observing positive and negative AGN feedback. Nature Astronomy 2, pp. 179–180. External Links: Document, 1802.10305, ADS entry Cited by: §1.
- GATOS: missing molecular gas in the outflow of NGC 5728 revealed by JWST. A&A 689, pp. A263. External Links: Document, 2406.17072, ADS entry Cited by: §1.